OpenStudy (anonymous):

A survey asked people which kind of pet they owned. The results are shown in the table below. Of the women who were surveyed, what is the relative frequency that owned cats?

OpenStudy (anonymous):

I need to see the table

OpenStudy (anonymous):

relative frequency is how often something happens dividied by all outcomes. For example : Lets say , if your team has won 10 games out of 20 games played, then the relative frequency is 10/20 = 1/2 = 50% I can't really do anything else for you since I have no table to look at